Let f(x)=x^3-8x^2+17x-9. Use the factor theorem to find other solutions to
f(x)-f(1)=0, besides x=1.

Answer by stanbon(75887)

Let f(x)=x^3-8x^2+17x-9. Use the factor theorem to find other solutions to
f(x)-f(1)=0, besides x=1.
f(1)=1-8+17-9=1
EQUATION:
f(x)-1=0
x^3-8x^2+17x-9-1=0
x^3-8x^2+17x-10=0
Using synthetic division find zeroes at x=1, or 2, or 5.
Therefore the factors are:
(x-1)(x-2)(x-5)=0
Cheers,
Stan H.
Stan H.